OpenStudy (anonymous):

If N = {all real numbers} and subset S = {multiples of 4 less than zero}, what is S′?

OpenStudy (anonymous):

What does the ' mean in this notation? Also, \(\mathbb{R}\) is used for all reals. Also, \[S:=\{4k: 4k<0 \wedge k \in \mathbb{Z}\}\]

OpenStudy (anonymous):

\[\mathbb{N}\] is usually reserved for natural numbers \[\{1,2,3,...\}\] just thought i would mention it

OpenStudy (anonymous):

actually the more i look, the more the problem is confusing me

OpenStudy (anonymous):

Is it one of these {multiples of 4 greater than zero} {even numbers less than zero} {odd numbers less than zero} {all real numbers excluding multiples of 4 less than zero}

OpenStudy (anonymous):

if we are living in the universe of real numbers, all can be written as multiples of 4

OpenStudy (anonymous):

Disregard my definition for \(S\).

OpenStudy (anonymous):

@Limitless am i being silly here? if the universe is \(\mathbb{R}\) does it make sense to talk about multiples of 4??

OpenStudy (anonymous):

I don't understand what \(S'\) means. @satellite73, I think you are correct about that. But, I do not see that as relevant because I don't know what's being asked.

OpenStudy (anonymous):

@CallMeKelly your teacher wants the last one as an answer

OpenStudy (anonymous):

this one "{all real numbers excluding multiples of 4 less than zero} "

OpenStudy (anonymous):

Is \(S'\) the elements not in \(S\)?

OpenStudy (anonymous):

i suppose i should leave my critique to myself so as not to confuse, but i cant help remarking that this is a really poor question on several levels

OpenStudy (zarkon):

S' is the complement of S

OpenStudy (anonymous):

@Zarkon am i being stupid here?

OpenStudy (zarkon):

no matter what the universal set is I take multiples of a natural number to be any integer times that number

OpenStudy (anonymous):

ok in any case the answer wanted is the last one

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

then my only objection is "let N be all real numbers" but whatever, i guess you can use any letter you like

OpenStudy (zarkon):

I object to that too...will lead to confusion down the road.

OpenStudy (anonymous):

This question doesn't seem well-motivated. If they're going to teach complements, why not use more important sets than just some arbitrarily defined set?

OpenStudy (zarkon):

whenever I teach this stuff I always start with simple finite sets

OpenStudy (anonymous):

I can understand that. I guess it really depends on your audience. Different people prefer different approaches to the same thing.

OpenStudy (anonymous):

i am still somewhat baffled by multiples of 4 in \(\mathbb{R}\) what is wrong with "integer multiples of 4"?

OpenStudy (zarkon):

that would be better wording